OpenStudy (anonymous):

Rewrite the expression using a radical. (a) 9 + x^(3/2) (b) (9 + x)^(3/2)

OpenStudy (saifoo.khan):

\[\Large 9 + x^\frac32 \to 9 + \sqrt[2]{x^3}\]

OpenStudy (mathstudent55):

\[(9 + x)^{\frac{ 3 }{ 2 }} = \sqrt{(9 + x)^{3}}\]

OpenStudy (anonymous):

I thought those were the correct answers also but when I type them in it says they are wrong.

OpenStudy (saifoo.khan):

Might be some input error?

OpenStudy (anonymous):

here's what I typed in.

OpenStudy (mathstudent55):

You are correct.